OpenStudy (anonymous):

How do you factor the difference of two squares? How do you factor the perfect square trinomial? How do you factor the sum and difference of two cubes? Which of these three makes the most sense to you? Explain why.

hero (hero):

Hint: Difference of Squares: \(a^2 - b^2 = (a+b)(a-b)\)

hero (hero):

Perfect square trinomial: \((a+b)^2 = a^2 + 2ab + b^2\)

hero (hero):

\(a^3 + b^3 = (a+b)(a^2 - ab + b^2)\) \(a^3-b^3 = (a-b)(a^2 + ab + b^3)\)

OpenStudy (mathmate):

The last term is b-squared! :)

hero (hero):

Yup, my bad.

OpenStudy (mathmate):

np